# Improving the Efficiency of Visually Augmented Language Models

### v1

@@ Abstract
Paula Ontalvilla HiTZ Center - Ixa, University of the Basque Country UPV/EHU paula.ontalvilla@ehu.eus Aitor Ormazabal HiTZ Center - Ixa, University of the Basque Country UPV/EHU aitor.ormazabal@ehu.eus Gorka Azkune HiTZ Center - Ixa, University of the Basque Country UPV/EHU gorka.azcune@ehu.eus
@@ 4 Experimental setup
Text Corpus. Following the original VaLM (Wang et al., 2022), we use the English corpus of CC-100 (Conneau et al., 2020) as the pre-training text corpus for all models. Due to limited access to compute, we only consume 10.5B tokens for pre-training, which accounts for around 19% of the English portion of the corpus.
Image data and retrieval module. VaLM requires an image database and a vector retrieval module. We use a FAISS (Johnson et al., 2019) index, trained on the exact same setup as the original VaLM, based on the GPT2-Small architecture (see Appendix A).
Pre-training hyperparameters. We train both models on the exact same setup, following a configuration similar to the original VaLM. Details can be found in Appendix A.
@@
Evaluation. We evaluate our models on VLU, NLU and LM tasks (see Appendix B for details).
@@
Scaling Blind-VaLM. It is important to note that, since Blind-VaLM does not require an actual image retrieval step, it is significantly more efficient at both training and inference time.222The efficiency gains are two-fold: 1) since we only use a single CLIP text representation in the Fusion Layer, this layer requires less floating-point operations (FLOPs), and 2) since we do away with the dense vector database retrieval, training and inference latency is significantly reduced, leading to improved wall-clock efficiency. Taking advantage of this increased efficiency, we explore two ways of scaling up Blind-VaLM, within the compute budget of our VaLM baseline.
Scaling model size. We train Blind-VaLM-Medium, by switching the LM backbone architecture to GPT2-Medium instead of GPT2-Small. See Appendix A for details. This larger model was trained on 595 GPU-hours, still within the compute budget of the VaLM baseline.
Scaling pre-training compute. We train Blind-VaLM+, by simply continuing pre-training the baseline Blind-VaLM for longer, until we reach a total of 88255 steps, which corresponds to 23.1B tokens (42% of CC-100). Again, thanks to the increased efficiency of our approach, this model is compute-matched333Note that we use compute-matched in the wall-clock time sense here, not the FLOPs sense, since the time consuming dense vector retrieval step we remove is not reflected by FLOPs. to the original VaLM baseline, taking a total of 1.17K GPU-hours to train.
@@ 5 Results
Blind-VaLM matches VaLM on VLU, NLU and LM tasks. Table 1 shows results for Blind-VaLM and VaLM trained on the same setup, as described in §4. We observe that our approach matches the original VaLM on VLU tasks, as well as NLU and LM tasks: it achieves an average score points higher in VLU tasks, and outperforms VaLM on 6/7 NLU & LM tasks. This supports our hypothesis that actually retrieving and encoding images is not required for visual augmentation, since simply utilizing textual representations from an already visually grounded text encoder works equally well. Additionally, as described in section §4 Blind-VaLM is 2.2x faster to train, since it skips the time consuming vector retrieval step. Speedups are even more significant at inference time, since generation is not as compute-bound and retrieval latency plays a bigger role.
Blind-VaLM outperforms VaLM, when trained within the same compute budget. Table 2 shows results for two scaled-up Blind-VaLM variants, obtained through scaling either pre-training compute or model size, as described in §4. We observe that both variants outperform VaLM, while being trained within the same compute budget as it. For example, in the case of Blind-VaLM-Medium, we outperform VaLM by points on average for VLU tasks, and outperform VaLM on 6/7 NLU & LM tasks.

### v2

@@ Abstract
Paula Ontalvilla Aitor Ormazabal Gorka Azkune HiTZ Center - Ixa, University of the Basque Country (UPV/EHU) {paula.ontalvilla, aitor.ormazabal, gorka.azcune}@ehu.eus
@@ 4 Experimental setup
Text Corpus.
Following the original VaLM (Wang et al., 2022), we use the English corpus of CC-100 (Conneau et al., 2020) as the pre-training text corpus for all models. Due to limited access to compute, we only consume 10.5B tokens for pre-training, which accounts for around 19% of the English portion of the corpus.
Image data and retrieval module.
Pre-training hyperparameters.
We train both models on the exact same setup, following a configuration similar to the original VaLM. Details can be found in Appendix A.
@@
Evaluation.
We evaluate our models on VLU, NLU and LM tasks (see Appendix B for details).
@@
Scaling Blind-VaLM.
It is important to note that, since Blind-VaLM does not require an actual image retrieval step, it is significantly more efficient at both training and inference time.222The efficiency gains are two-fold: 1) since we only use a single CLIP text representation in the Fusion Layer, this layer requires less floating-point operations (FLOPs), and 2) since we do away with the dense vector database retrieval, training and inference latency is significantly reduced, leading to improved wall-clock efficiency. Taking advantage of this increased efficiency, we explore two ways of scaling up Blind-VaLM, within the compute budget of our VaLM baseline.
(i) Scaling model size.
We train Blind-VaLM-Medium, by switching the LM backbone architecture to GPT2-Medium instead of GPT2-Small. See Appendix A for details. This larger model was trained on 595 GPU-hours, still within the compute budget of the VaLM baseline.
(ii) Scaling pre-training compute.
We train Blind-VaLM+, by simply continuing pre-training the baseline Blind-VaLM for longer, until we reach a total of 88255 steps, which corresponds to 23.1B tokens (42% of CC-100). Again, thanks to the increased efficiency of our approach, this model is compute-matched333Note that we use compute-matched in the wall-clock time sense here, not the FLOPs sense, since the time consuming dense vector retrieval step we remove is not reflected by FLOPs. to the original VaLM baseline, taking a total of 1.17K GPU-hours to train.
@@ 5 Results
Blind-VaLM matches VaLM on VLU, NLU and LM tasks.
Table 1 shows results for Blind-VaLM and VaLM trained on the same setup, as described in §4. We observe that our approach matches the original VaLM on VLU tasks, as well as NLU and LM tasks: it achieves an average score points higher in VLU tasks, and outperforms VaLM on 6/7 NLU & LM tasks. This supports our hypothesis that actually retrieving and encoding images is not required for visual augmentation, since simply utilizing textual representations from an already visually grounded text encoder works equally well. Additionally, as described in section §4 Blind-VaLM is 2.2x faster to train, since it skips the time consuming vector retrieval step. Speedups are even more significant at inference time, since generation is not as compute-bound and retrieval latency plays a bigger role.
Blind-VaLM outperforms VaLM, when trained within the same compute budget.
Table 2 shows results for two scaled-up Blind-VaLM variants, obtained through scaling either pre-training compute or model size, as described in §4. We observe that both variants outperform VaLM, while being trained within the same compute budget as it. For example, in the case of Blind-VaLM-Medium, we outperform VaLM by points on average for VLU tasks, and outperform VaLM on 6/7 NLU & LM tasks.
@@ 7 Limitations
Acknowledgments
This work is partially supported by the Ministry of Science and Innovation of the Spanish Government (AWARE project TED2021-131617B-I00, DeepR3 project TED2021-130295B-C31), project funded by MCIN/AEI/10.13039/501100011033 and European Union NextGeneration EU/PRTR, and the Basque Government (IXA excellence research group IT1570-22 and IKER-GAITU project).